# One missing piece in Vision and Language: A Survey on Comics Understanding

### v1

@@ Abstract
Vision-language models have recently evolved into versatile systems capable of high performance across a range of tasks, such as document understanding, visual question answering, and grounding, often in zero-shot settings. Comics Understanding, a complex and multifaceted field, stands to greatly benefit from these advances. Comics, as a medium, combine rich visual and textual narratives, challenging AI models with tasks that span image classification, object detection, instance segmentation, and deeper narrative comprehension through sequential panels. However, the unique structure of comics—characterized by creative variations in style, reading order, and non-linear storytelling—presents a set of challenges distinct from those in other visual-language domains. In this survey, we present a comprehensive review of Comics Understanding  from both dataset and task perspectives. Our contributions are fivefold: (1) We analyze the structure of the comics medium, detailing its distinctive compositional elements; (2) We survey the widely used datasets and tasks in comics research, emphasizing their role in advancing the field; (3) We introduce the Layer of Comics Understanding  (LoCU) framework, a novel taxonomy that redefines vision-language tasks within comics and lays the foundation for future work; (4) We provide a detailed review and categorization of existing methods following the LoCU framework; (5) Finally, we highlight current research challenges and propose directions for future exploration, particularly in the context of vision-language models applied to comics. This survey is the first to propose a task-oriented framework for comics intelligence and aims to guide future research by addressing critical gaps in data availability and task definition. A project associated with this survey is available at https://github.com/emanuelevivoli/awesome-comics-understanding.
Index Terms:
@@
Survey, Comics, Manga, Layers of Understanding, Comics Understanding, Vision-Language tasks
1 Introduction
@@
Comics serve as a sophisticated medium that combines visual and textual elements to tell stories [1]. Applying recent machine learning tools to understand comics is complicated. This complexity comes from its unique domain that, driven by the author’s creativity, exhibits unique variations in both style and content. For example, modeling the drawing style is complicated, the reading order changes from one book to another, and, unlike natural images, comics frequently depict transformations that defy physical laws, adding another layer of complexity.
@@
Motivated by these challenges, advanced artificial intelligence (AI) approaches have been increasingly applied to Comics Understanding. Given that AI thrives on tackling complex and diverse tasks, many researchers are now focusing on problems like object detection [2], semantic segmentation [3], optical character recognition (OCR) [4], recurrence of characters and objects in varying contexts [5], pose estimation [6], depth estimation [7], and integration of visual and textual narratives [8]. Comics also include complex narrative elements such as satire and irony and spatial and temporal dynamics within the story that have not been deeply investigated. Many of these tasks, individually or in combination, represent the cutting edge of advanced AI, particularly in the area of Multimodal Vision-Language understanding [9, 10, 11, 12, 13].
@@
Figure 1: Number of publications on Comics and Manga (from Google Scholar). The publications have been filtered by keywords (manga, comics, graphic novels), topics and journals (computer science).
@@
Historically, in machine learning (and deep learning), innovation has been driven by the available data, both in terms of tasks and more capable models. As new types of annotated data and more complex tasks emerged, the need for innovative approaches grew, particularly in domains that combine multiple modalities. The domain of multimodal learning, especially the convergence of vision and language, has witnessed significant advancements. These advancements include the shift away from independent uni-modal representations [14, 15] toward the alignment and integration of different modalities [16, 17, 18, 19, 20, 21] and the creation of multimodal embedding space combining aligned uni-modal inputs [22, 23], with early and late fusion. Trained on extensive multimodal datasets, these models have shown remarkable abilities in tasks like multimodal retrieval [14, 22, 23], visual question answering, image captioning [17], reasoning [20, 21], and retrieval-augmented multimodal generation [24] as well as image generation [25, 26].
The comics domain is uniquely well-suited to driving advancements in these types of multimodal reasoning models. Research in comics has extensively explored a range of questions, from the human ability to derive meaning from sequential images [27] to machine interpretation of comic strips, particularly through closure tasks [8].
@@
With the continuous advancement in AI models and the intensive interest in harvesting vast knowledge from Comics, a lot of work is being published constantly. This is evidenced by a great number of recent papers as shown in Fig. 1. These contributions shifted the landscape around Comics Understanding, leaving previous surveys outdated [6]. Nowadays, the Comics Understanding  research community lacks a comprehensive survey that catalogs these papers and the various challenges in the Vision-Language domain.
@@
To address this need, we introduce in this paper the novel Layer of Comics Understanding (LoCU) framework. This taxonomy, illustrated in Table I, delineates tasks by their input/output modalities and the spatio-temporal dimensions necessary for reasoning over comic data, inviting a structured and progressive approach to comics analysis. Motivated by the uniquely broad task distribution of the comics medium, this survey aggregates and examines all extant research on machine learning in comics, aimed at highlighting critical unresolved issues in the field. It is the first to review the breadth of comics research in a hierarchical structure, and to our knowledge, the first to propose a big-picture taxonomy of comics intelligence from a task-oriented perspective.
@@
Figure 2: Visualization of the Layers of Comics Understanding tasks. From panel level to multipage, from unimodal to multimodal, from simplest to more complex.
@@
The rest of this survey is organized as follows. Section 2 introduces the Comics medium as a composition of Visual-Language tasks, and several related surveys. Section 3 describes the Foundations of Comics, including their structure, similarities, and differences among various types. Section 4 introduces existing tasks and the commonly used datasets in Comics Understanding, also focusing on evaluation procedures. In Section 5 we propose the Layer of Comics Understanding taxonomy (LoCU), and for each layer a collection of tasks grouped by difficulty (see Annex LABEL:annex:values). In Section 6 we start with Layer 1: Tagging and Augmentation, Section 7 with Layer 2: Grounding, Analysis, and Segmentation, in Section 8 with Layer 3: Retrieval and Modification, Section 9 with Layer 4: Understanding, and in Section 10 with Layer 5: Generation and Synthesis. Finally, we conclude with Section 11.
2 Background
@@
2.1 The Comics Research Epochs
@@
2.1.1 Traditional Machine Learning
Early machine learning approaches heavily depended on feature engineering, utilizing hand-crafted features [28, 29, 30], which were capable of, e.g., classify the style of a manga using SVMs [31], retrieve a manga character from a sketch with edge orientation histograms [32] or segment comic balloons with energy functions and active contour models [33]. However, these paradigms require domain experts to craft effective features for specific tasks, which do not cope well with complex tasks and have poor generalizability. Moreover, the feature engineering obstacles the extension of the Comics Understanding to more tasks.
2.1.2 Deep Learning approaches
With the development of deep learning, researchers have achieved major improvements by using end-to-end trainable deep neural networks (DNNs). These models eliminate the need for complex feature engineering and shift the focus to designing network architectures that learn features automatically. Applying these paradigms to comics came as a natural step. Models became better at detection [34, 35] and segmentation [36, 37, 38], and more complex tasks started being investigated and solved (e.g. visual and textual closure [8]). However, the turn from traditional machine learning toward deep learning raises a new grand challenge: the laborious collection of large-scale, task-specific, and crowd-labeled data in DNN training. Moreover, while deep learning methods excel in specific tasks, they are often difficult to adapt to multi-task. In the domain of comics, where data is scarce, protected by copyright, and challenging to annotate, these constraints have significantly hindered the development of more generalized, multi-task models for Comics Understanding.
2.1.3 Modern Foundational models
The era of foundational models marks a significant evolution in the application of AI to comics, with models that can learn from vast amounts of unstructured data to perform a variety of complex tasks [14, 39, 40]. These models relax previous limitations by “solving” simple yet important tasks as a zero-shot counter effect, e.g. panel, character, and text-box detection [41]. By leveraging these foundational models, more advanced tasks can be proposed and takled [42, 43], opening the doors to yet unexplored more complex ones.
TABLE I: Overview of Vision-Language Tasks of the Layers of Comics Understanding. The ranking is based on input and output modalities and dimensions, as illustrated in Appendix B. Taxonomy.
@@ 2.2 Relevant Surveys
To the best of our knowledge, this is the first survey that reviews comics papers with a task-oriented approach for Vision-Language. Several relevant surveys have been conducted, examining broader aspects of visual art [44] and, more specifically, comics [45, 6, 46]. Notably, Augereau et al. [6] categorized comic research into three primary areas: (i) content analysis, (ii) content generation and adaptation, and (iii) user interaction, which remain relevant entry points for understanding this field’s research and applications. Since those surveys were published, foundational models have emerged, enabling significant advancements, particularly due to their ability to generalize across tasks without task-specific training. These models have unlocked new possibilities in comics research, allowing automatic annotation and analysis, among other tasks.
@@
Building on these advancements, our survey focuses on two key aspects: 1) the recent progress in Comics Understanding tasks and datasets, and 2) the application of Vision-Language tasks to the unique characteristics of comics.
3 Comics Foundations
@@
Our preliminary analysis prioritizes understanding the structural intricacies of comics. The term “comics” broadly denotes the entire medium in its uncountable form, while “a comic” refers to a singular entity within this medium, such as a comic book or strip. The structure of comic books has been a subject of extensive study, examining aspects like their commonalities, layout structures [47], content arrangement, narrative transitions [48, 27, 49, 50, 51], and stylistic narrative variations [52, 53, 54].
@@
Figure 3: Anatomy of Comic Page Elements: This illustration delineates the various components found on comic pages including object detection, linkages between elements, the designated reading order of texts and panels, as well as key textual features like character names—all arranged in a logical sequence throughout the panels and pages.
@@
Typically, comics are presented in single or double-page formats, influenced by the artist’s style. They narrate stories across various genres or series, such as superheroes and science fiction tales, sometimes interlinking through crossovers. Notably, artistic consistency within a series or genre, often achieved through collaborative efforts, is a key feature. Each comic page (as illustrated in Figure 3) encapsulates individual scenes within panels of diverse types - from splash panels for opening scenes to widescreen panels for expansive narratives. These panels convey interactions and dialogues through speech balloons, thought bubbles, and expressive onomatopoeia, enriching the narrative experience. Artistic variations in style and color also play a significant role in comics, ranging from detailed realism to abstract forms, with the color palette setting the mood and atmosphere.
Despite rigorous analyses [55], no significant differences have been identified in terms of attention, subjectivity, and viewpoint between Manga and American comics. Hence, these distinctions are better explored within their distinct historical, cultural, and artistic contexts:
@@
Manga: Emerging in post-World War II Japan and influenced by diverse sources, including Japanese traditional art and American comics, manga is recognized for its unique black-and-white style, though colored variants exist. Covering a broad range of genres, Manga is read right-to-left, top-to-bottom. This style is marked by a distinctive “Japanese Visual Language” that is consistent across genres [56].
Comic Formalisms. Comics are defined as sequential visual narratives that utilize a combination of images and text within distinct frames (panels) to convey a story or information. Each comic, or manga, can be represented as a document $D$. Each document $D$ consists of a series of pages $P=\{p_{1},p_{2},\ldots,p_{n}\}$. A page $p_{i}$ can be described by its layout and the elements it contains: (i) a spatial arrangement of panels $\{c_{1},c_{2},\ldots,c_{m}\}$, also known as Layout; (ii) each panel $c_{i}$ is a bounded area containing images and possibly text; panels are the primary unit for storytelling in comics; (iii) gutters: the space between panels, often used by artist to leave readers imagining the missing piece of the story, which helps in sequencing and timing; and (iv) text elements, including balloons (speech or thought), captions, and sound effects (onomatopoeias). Each panel $c_{i}$ contains characters (depicted through various visual styles and expressions), backgrounds (visual scenery behind the main characters or action), and foregrounds (objects and elements in front of the characters).
4 Tasks and Datasets
@@
This section summarizes the commonly used Comics datasets, as detailed in Table II. Some of these datasets are either intended for training or evaluation, depending on their sizes.
TABLE II: Overview of Comic/Manga Datasets and Tasks, their information (availability, published year, source), and properties (languages, number of comic/manga books and pages). The table rows are repeated according to the supported Tasks. Accessibility is indicated with for no longer existing datasets, indicates existing but not accessible, and means existing and accessible. The link [proj] directs to the project websites, while [data] directs to dataset websites. For CoMix, mix means that it inherits from a mixture of four datasets [57, 8, 58, 41]. For VLRC, Analysis indicates annotations are about statistical analysis of panel size, color, layout, etc. However, the dataset does not provide images and annotations have been gathered directly on physical books.
Task From Name Year Access. Language Origin # books # pages
Image Classification Sequencity [59] 2017 EN, JP - - 140000
BAM! [60] 2017 - - - 2500000
[61] Manga109 [61, 35] [proj][data] 2018 JP 1970-2010 109 21142
[8] EmoRecCom [62] [proj] [data] 2021 EN 1938-1954 - -
Object Detection Fahad18 [63] [proj] 2012 - - - 586
eBDtheque [57] [proj][data] 2013 EN, FR, JP 1905-2012 25 100
[57] sun70 [64] 2013 FR - 6 60
COMICS [8] [proj] [data] 2017 EN 1938-1954 3948 198657
BAM! [60] 2017 - - - 2500000
[61] JC2463 [34] 2017 JP - 14 2463
AEC912 [34] 2017 EN, FR - - 912
GCN [65] [proj] [data] 2017 EN, JP 1978-2013 253 38000
[59] Sequencity612 [59] 2017 EN, JP - - 612
[57] SSGCI [66] [proj] [data] 2016 EN, FR, JP 1905-2012 - 500
[61] Comics3w [67] [proj] 2017 JP, EN - 103 29845
comics2k [68] [proj] [data] 2018 - - - -
DCM772 [58] [proj] [data] 2018 EN 1938-1954 27 772
[61] Manga109 [61, 35] [proj][data] 2018 JP 1970-2010 109 21142
BCBId [69] [proj] [data] 2022 BN - 64 3327
[61] COO [4] [proj] [data] 2022 JP 1970-2010 109 10602
[8] COMICS-Text+ [70] [proj] [data] 2022 EN 1938-1954 3948 198657
PopManga [41] [proj] [data] 2024 EN 1990-2020 25 1925
mix CoMix [42] [proj] [data] 2024 EN, FR 1938-2023 100 3800
Re-Identification Fahad18 [63] [proj] 2012 - - - 586
Ho42 [71] 2013 - - - 42
[61] Manga109 [61, 35] [proj][data] 2018 JP 1970-2010 109 21142
PopManga [41] [proj] [data] 2024 EN 1990-2020 25 1925
mix CoMix [42] [proj] [data] 2024 EN, FR 1938-2023 100 3800
Linking eBDtheque [57] [proj][data] 2013 EN, FR, JP 1905-2012 25 100
[57] sun70 [64] 2013 FR - 6 60
GCN [65] [proj] [data] 2017 EN, JP 1978-2013 253 38000
[61] Manga109 [61, 35] [proj][data] 2018 JP 1970-2010 109 21142
PopManga [41] [proj] [data] 2024 EN 1990-2020 25 1925
mix CoMix [42] [proj] [data] 2024 EN, FR 1938-2023 100 3800
Segmentation [59] Sequencity4k [38] 2020 EN, FR, JP - - 4479
Dialog generation PopManga [41] [proj] [data] 2024 EN 1990-2020 25 1925
mix CoMix [42] [proj] [data] 2024 EN, FR 1938-2023 100 3800
Analysis VLRC [47] [proj] [data] 2023 JP, FR, EN, 6+ 1940-present 376 7773
4.1 Annotations overview

### v2

@@ Abstract.
Vision-language models have recently evolved into versatile systems capable of high performance across various tasks, often in zero-shot settings. Comics Understanding, a complex and multifaceted field, stands to benefit from these advances greatly. As a medium, comics combine rich visual and textual narratives, challenging AI models with tasks that span image classification, object detection, instance segmentation, and deeper narrative comprehension through sequential panels. However, the unique structure of comics—marked by stylistic variation, reading order complexity, and non-linear storytelling—poses distinct challenges. In this survey, we provide a comprehensive review of Comics Understanding from both dataset and task perspectives. Our contributions include: (1) analyzing the unique structure and elements of the comics medium; (2) surveying key datasets and tasks in comics research; (3) introducing the Layer of Comics Understanding (LoCU) framework, a novel taxonomy for redefining vision-language tasks in comics; (4) categorizing existing methods using the LoCU framework; and (5) identifying research challenges and future directions for applying vision-language models to comics. This survey pioneers a task-oriented framework for comics intelligence, aiming to guide future research by addressing gaps in data and task definition. A project associated with this survey is available at https://github.com/emanuelevivoli/awesome-comics-understanding.
@@
Comics understanding, Vision and Language tasks, Comics, Manga, Sequential Visual Art
$c$s: Computing methodologies Information extractionccs: Computing methodologies Computer vision representationsccs: Computing methodologies Image representationsccs: Computing methodologies Computer graphicsccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processing
1. Introduction
@@
Comics serve as a sophisticated medium that combines visual and textual elements to tell stories (Scholarsarchive et al., 2018). Applying recent machine learning tools to understand comics is complicated. This complexity comes from its unique domain that, driven by the author’s creativity, exhibits unique variations in both style and content. For example, modeling the drawing style is complicated, the reading order changes from one book to another, and, unlike natural images, comics frequently depict transformations that defy physical laws, adding another layer of complexity.
@@
Motivated by these challenges, advanced artificial intelligence (AI) approaches have been increasingly applied to Comics Understanding. Given that AI thrives on tackling complex and diverse tasks, many researchers are now focusing on problems like object detection (Vivoli et al., 2024c), semantic segmentation (Bhattacharjee and Aydemir, 2024), optical character recognition (OCR) (Baek et al., 2022), recurrence of characters and objects in varying contexts (Li et al., 2021), pose estimation (Augereau et al., 2018), depth estimation (Bhattacharjee et al., 2021), and integration of visual and textual narratives (Iyyer et al., 2016). Comics also include complex narrative elements such as satire and irony and spatial and temporal dynamics within the story that have not been deeply investigated. Many of these tasks, individually or in combination, represent the cutting edge of advanced AI, particularly in the area of Multimodal Vision-Language understanding (Liu et al., 2023a; Li et al., 2024b; Xue et al., 2024; Laurençon et al., 2024; Yao et al., 2024).
@@
Figure 1. Number of publications on Comics and Manga (from Google Scholar). The publications have been filtered by keywords (manga, comics, graphic novels), topics, and journals (computer science).
@@
Historically, in machine learning (and deep learning), innovation has been driven by the available data, both in terms of tasks and more capable models. As new types of annotated data and more complex tasks emerged, the need for innovative approaches grew, particularly in domains that combine multiple modalities. The domain of multimodal learning, especially the convergence of vision and language, has witnessed significant advancements. These advancements include the shift away from independent uni-modal representations (Radford et al., 2021; Jia et al., 2021) toward the alignment and integration of different modalities (Cho et al., 2021; Li et al., 2023b; Singh et al., 2021; Zhu et al., 2023a; Chen et al., 2023b; Wang et al., 2023b) and the creation of multimodal embedding space combining aligned uni-modal inputs (Girdhar et al., 2023; Zhu et al., 2023b), with early and late fusion. Trained on extensive multimodal datasets, these models have shown remarkable abilities in tasks like multimodal retrieval (Radford et al., 2021; Girdhar et al., 2023; Zhu et al., 2023b), visual question answering, image captioning (Li et al., 2023b), reasoning (Chen et al., 2023b; Wang et al., 2023b), and retrieval-augmented multimodal generation (Yasunaga et al., 2023) as well as image generation (Gal et al., 2022; Yang et al., 2022).
@@
The comics domain is uniquely well-suited to driving advancements in these types of multimodal reasoning models. Research in comics has extensively explored a range of questions, from the human ability to derive meaning from sequential images (Cohn, 2013b) to machine interpretation of comic strips, particularly through closure tasks (Iyyer et al., 2016). With the continuous advancement in AI models and the intensive interest in harvesting vast knowledge from Comics, a lot of work is being published constantly. This is evidenced by a great number of recent papers as shown in Fig. 1. These contributions shifted the landscape around Comics Understanding, leaving previous surveys outdated (Augereau et al., 2018). Nowadays, the Comics Understanding  research community lacks a comprehensive survey that catalogs these papers and the various challenges in the Vision-Language domain.
@@
To address this need, we introduce in this paper the novel Layer of Comics Understanding (LoCU) framework. This taxonomy, illustrated in Table 1, delineates tasks by their input/output modalities and the spatio-temporal dimensions necessary for reasoning over comic data, inviting a structured and progressive approach to comics analysis. Motivated by the uniquely broad task distribution of the comics medium, this survey aggregates and examines all extant research on machine learning in comics, aimed at highlighting critical unresolved issues in the field. It is the first to review the breadth of comics research in a hierarchical structure, and to our knowledge, the first to propose a big-picture taxonomy of comics intelligence from a task-oriented perspective.
@@
Figure 2. Visualization of the tasks in Layers of Comics Understanding. From panel level to multipage, from unimodal to multimodal, and from simplest to more complex.
@@
The rest of this survey is organized as follows. Section 2 introduces the Comics medium as a composition of Visual-Language tasks, and several related surveys. Section 3 describes the Foundations of Comics, including their structure, similarities, and differences among various types. Section 4 introduces existing tasks and the commonly used datasets in Comics Understanding, also focusing on evaluation procedures. In Section 5 we propose the Layer of Comics Understanding taxonomy (LoCU), and for each layer a collection of tasks grouped by difficulty. In Section 6 we start with Layer 1: Tagging and Augmentation, Section 7 with Layer 2: Grounding, Analysis, and Segmentation, in Section 8 with Layer 3: Retrieval and Modification, Section 9 with Layer 4: Understanding, and in Section 10 with Layer 5: Generation and Synthesis. Finally, we conclude with Section 11.
2. Background
@@
2.1. The Comics Research Epochs
@@
Traditional Machine Learning. Early machine learning approaches heavily depended on feature engineering, utilizing hand-crafted features (Chou and Shih, 2011; Amirshahi et al., 2012; Sun and Kise, 2013), which were capable of, e.g., classify the style of a manga using SVMs (Chu and Chao, 2014), retrieve a manga character from a sketch with edge orientation histograms (Matsui et al., 2014) or segment comic balloons with energy functions and active contour models (Rigaud et al., 2013). However, these paradigms require domain experts to craft effective features for specific tasks, which do not cope well with complex tasks and have poor generalizability. Moreover, the feature engineering obstacles the extension of the Comics Understanding to more tasks.
Deep Learning approaches. With the development of deep learning, researchers have achieved major improvements by using end-to-end trainable deep neural networks (DNNs). These models eliminate the need for complex feature engineering and shift the focus to designing network architectures that learn features automatically. Applying these paradigms to comics came as a natural step. Models became better at detection (Qin et al., 2017; Ogawa et al., 2018) and segmentation (Dubray and Laubrock, 2019; Dutta et al., 2021; Nguyen et al., 2020), and more complex tasks started being investigated and solved (e.g. visual and textual closure (Iyyer et al., 2016)). However, the turn from traditional machine learning toward deep learning raises a new grand challenge: the laborious collection of large-scale, task-specific, and crowd-labeled data in DNN training. Moreover, while deep learning methods excel in specific tasks, they are often difficult to adapt to multi-task. In the domain of comics, where data is scarce, protected by copyright, and challenging to annotate, these constraints have significantly hindered the development of more generalized, multi-task models for Comics Understanding.
Modern Foundational models. The era of foundational models marks a significant evolution in the application of AI to comics, with models that can learn from vast amounts of unstructured data to perform a variety of complex tasks (Radford et al., 2021; Oquab et al., 2023; Liu et al., 2023c). These models relax previous limitations by “solving” simple yet important tasks as a zero-shot counter effect, e.g. panel, character, and text-box detection (Sachdeva and Zisserman, 2024). By leveraging these foundational models, more advanced tasks can be proposed and takled (Vivoli et al., 2024a; Sachdeva et al., 2024), opening the doors to yet unexplored more complex ones.
Table 1. Overview of Vision-Language Tasks of the Layers of Comics Understanding. The ranking is based on input and output modalities and dimensions, as illustrated in Supplementary Materials.
@@ 2.2. Relevant Surveys
To the best of our knowledge, this is the first survey that reviews comics understanding publications with a task-oriented approach for Vision-Language. Several relevant surveys have been conducted, examining broader aspects of visual art (Santos et al., 2021) and, more specifically, comics (Augereau et al., 2017, 2018; Nairat, 2021). Notably, Augereau et al. (Augereau et al., 2018) categorized comic research into three primary areas: (i) content analysis, (ii) content generation and adaptation, and (iii) user interaction, which remain relevant entry points for understanding this field’s research and applications.
@@
In contrast, building on recent advancements, our survey focuses on two key aspects: 1) the recent progress in Comics Understanding tasks and datasets, and 2) the application of Vision-Language tasks to the unique characteristics of comics.
3. Comics Foundations
@@
Our preliminary analysis prioritizes understanding the structural intricacies of comics. The term “comics” broadly denotes the entire medium in its uncountable form, while “a comic” refers to a singular entity within this medium, such as a comic book or strip. The structure of comic books has been a subject of extensive study, examining aspects like their commonalities, layout structures (Cohn, 2023), content arrangement, narrative transitions (Cohn, 2013a, b; Cohn et al., 2014; Cohn, 2014, 2016), and stylistic narrative variations (Cohn et al., 2017; Ang, 2021; Klomberg et al., 2022).
@@
Figure 3. Anatomy of Comic Page Elements: This illustration delineates the various components found on comic pages including object detection, linkages between elements, the designated reading order of texts and panels, as well as key textual features like character names—all arranged in a logical sequence throughout the panels and pages.
@@
Typically, comics are presented in single or double-page formats, influenced by the artist’s style. They narrate stories across various genres or series, such as superheroes and science fiction tales, sometimes interlinking through crossovers. Notably, artistic consistency within a series or genre, often achieved through collaborative efforts, is a key feature. Each comic page (as illustrated in Figure 3) encapsulates individual scenes within panels of diverse types - from splash panels for opening scenes to widescreen panels for expansive narratives. These panels convey interactions and dialogues through speech or thought balloons, and expressive onomatopoeia, enriching the narrative experience. Artistic variations in style and color also play a significant role in comics, ranging from detailed realism to abstract forms, with the color palette setting the mood and atmosphere. Despite rigorous analyses (Cohn, 2011), no significant differences have been identified in terms of attention, subjectivity, and viewpoint between Manga and American comics. Hence, these distinctions are better explored within their distinct historical, cultural, and artistic contexts:
@@
Manga: Emerging in post-World War II Japan and influenced by diverse sources, including Japanese traditional art and American comics, manga is recognized for its unique black-and-white style, though colored variants exist. Covering a broad range of genres, Manga is read right-to-left, top-to-bottom. This style is marked by a distinctive “Japanese Visual Language” that is consistent across genres (Cohn and Ehly, 2016).
4. Tasks and Datasets
@@
This section summarizes the commonly used Comics datasets, as detailed in Table 2. Some of these datasets are either intended for training or evaluation, depending on their sizes.
[…]
From Name Year Acc. Lang. Origin # books # pages IC OD RI L S DG A
[…]
(Guérin et al., 2013) sun70 (Sun et al., 2013) 2013 FR - 6 60 x x
Ho42 (Ho et al., 2013) 2013 - - - 42 x
[…]
Sequencity (Nguyen et al., 2017) 2017 GB, JP - - 140000 x
BAM! (Wilber et al., 2017) 2017 - - - 2500000 x x
[…]
(Fujimoto et al., 2016) JC2463 (Qin et al., 2017) 2017 JP - 14 2463 x
AEC912 (Qin et al., 2017) 2017 GB, FR - - 912 x
[…]
(Nguyen et al., 2017) Sequencity612 (Nguyen et al., 2017) 2017 GB, JP - - 612 x
[…]
(Nguyen et al., 2017) Sequencity4k (Nguyen et al., 2020) 2020 GB, FR, JP - - 4479 x
[…]
4.1. Annotations overview